OpenStudy (love_to_love_you):

Please help! Will give medal and fan! Please. 1. Simplify the following expression and justify each step in your simplification. 5(7-x)+6x(3x+4) 2. A laptop advertisement states that internal parts operate at 55 degrees Celsius plus or minus 15 degrees. Write an absolute value equation to represent the situation. Then solve the equation and write it as a compound inequality. State the meaning of the solution. 3. Write an equation of a line that goes through point (4, -5) and is parallel to line y= 2x+8. Show all work.

OpenStudy (love_to_love_you):

4. Evaluate the expression for the given value of the variable. |-4b-8| + |-1-b^2| + 2b^3; b=-2 5. Write the expression in simplest form. 3(a^2+b) - 4(a^2-b) 6. Solve the equation. Check for extraneous solutions. 9|9-8x| = 2x+3 Please show work on all questions.

OpenStudy (anonymous):

hmmm one suggestion here :)) post one question at a time. haha but I will go ahead and solve #1 for ya :) 1. Simplify the following expression and justify each step in your simplification. 5(7-x)+6x(3x+4) so that will be 35-5x+18x^2+24x when u distribute. now combine like terms. 18x^2+19x+35 would be your final answer if you are just simplifying.

OpenStudy (love_to_love_you):

Thank you! Is that all the steps? That's it?

OpenStudy (love_to_love_you):

@Melodysim

OpenStudy (love_to_love_you):

Could you be amazing and help me with any of the others? :) @Melodysim

OpenStudy (anonymous):

yes and I can work #3 and #5 out for you. I was called in other question sorry

OpenStudy (anonymous):

kk #3 3. Write an equation of a line that goes through point (4, -5) and is parallel to line y= 2x+8. Show all work. so. parallel slopes are equal; 2 therefore, if you plug it in it will be -5=2(4)+b than you solve it for b. -5=8+b -13=b noww plug it back in to y=mx+b it will be -5=2x-13

OpenStudy (anonymous):

oops I meant y=2x-13

OpenStudy (love_to_love_you):

Ok, so instead of -5=2x-13 its y=

OpenStudy (anonymous):

yes :))

OpenStudy (anonymous):

okay than #5 3(a^2+b) - 4(a^2-b) distribute!! 3a^2+3b-4a^2+4b now you combine like terms -a^2+7b this would be your simplest form!!

OpenStudy (love_to_love_you):

Thank you sooooooo much! You're a lifesaver. Really :) Now I have 2, 4, and 6. Do you recommend posting separately or together? @Melodysim

OpenStudy (anonymous):

separately. because ppl tend to freak out when they see problems all together haha like I kinda did :)) @love_to_love_you

OpenStudy (love_to_love_you):

Haha alrighty. Thanks :)

OpenStudy (anonymous):

Blah. no problem!